The measure of one angle of a triangle is 28 more than the measure of the smallest angle of the triangle. The measure of the third angle is twice the measure of the smallest angle. Find all measures.
Answer by rfer(10417)

x+x+28+2x=180
4x+28=180
4x=152
x=38 degrees
x+28=66 degrees
2x=76 degrees